OpenStudy (anonymous):

Please Help!!!! Fan and Medals!!!!! A bag contains 6 blue marbles, 3 green marbles, and 2 yellow marbles. Ava will select a marble from the bag without looking. What is the probability that she will select a green or a yellow marble?

OpenStudy (anonymous):

\[A. \frac{ 6 }{ 121 }\] \[B. \frac{ 5 }{ 22 }\] \[C. \frac{ 5 }{ }\] \[D. \frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

The probability of choosing a green marble is \[3\over11\] The probability of choosing a yellow marble is \[2\over11\] So the answer should be the combined fractions: \[\frac{ 3 }{ 11 } + \frac{ 2 }{ 11 }\]